OpenStudy (anonymous):

2. The sum of the measures of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X? (1 point) A.20° B.35° C.50° D.65°

OpenStudy (anonymous):

Angle Y=y Angle X=2y-30 y+2y-30=90 3y-30=90 3y=120 y=40 x=2y-30 x=2(40)-30 x=50